Exercise 22-22

On January 1, 2017, Headland Co. purchased 22,000 shares (a 10% interest) in Elton John Corp. for $1,480,000. At the time, the book value and the fair value of John’s net assets were $12,100,000.

On July 1, 2018, Headland paid $3,340,000 for 44,000 additional shares of John common stock, which represented a 20% investment in John. The fair value of John’s identifiable assets net of liabilities was equal to their carrying amount of $13,200,000. As a result of this transaction, Headland owns 30% of John and can exercise significant influence over John’s operating and financial policies. (Any excess fair value is attributed to goodwill.)

John reported the following net income and declared and paid the following dividends.

Net Income

Dividend per Share

Year ended 12/31/17 $630,000 None
Six months ended 6/30/18 490,000 None
Six months ended 12/31/18 754,000 $1.50


Determine the ending balance that Headland Co. should report as its investment in John Corp. at the end of 2018.

Investment in Elton John Corp. $
Answer #1

Date

Particulars

Debit ($)

Credit ($)

01.01.2017

Investment in Elton John

1,480,000

       Cash

1,480,000

(To record investment)

01.07.2018

Investment in Elton John

3,340,000

        Cash

3,340,000

(To record investment)

In the year 2017, the dividend has not been paid and net income is 690,000.

The share of Headland would be $ 690,000 = $ 69,000

Hence, total carrying amount = 69000 +1480,000 =$ 1,549,000

For half of 2018, net income is $ 450,000, the share of Headland = 10% 0f 490,000 = $ 49,000

Total carrying amount = 1549,000 + 49,000 = $ 1,598,000

Now after July 2018,

Total investment in Elton john would be = 1598,000 + 3,340,000 =$ 4,938,000

Share in net income would be 30% of 754,000 = $ 226,200

Dividend received = 1.50 X 22000 X 3 = $ 99,000

It ought to be adjusted in carrying amount

Adjusted carrying amount = 4938,000 + 226,200 – 99000 = $ 5065,200

Hence, current investment in Elton John would be $ 5065,200.

Note that on 31/12/2018, 30% interest is there instead of 10%( 22,000 shares). So, 22000 has been multiplied by 3 to reflect current 30% interest share.

kindly upvote